# The Geometry of the Pivot: A Note on Lazy Pivoted Cholesky and Farthest Point Sampling

### v2

@@ 3.2 The Update Step as Orthogonal Projection
Note that by definition of orthogonal projection, the residual vector $r_{m}$ is orthogonal to the subspace $S_{m-1}$. Since the diagonal element serves as the normalization factor $L_{mm}=\sqrt{d[\pi_{m}]}=\|r_{m}\|$, the update step explicitly computes the projection onto the new normalized basis vector $e_{m}=r_{m}/\|r_{m}\|$:
$L_{i,m}=\frac{\langle\phi(x_{i}),r_{m}\rangle}{\|r_{m}\|}=\langle\phi(x_{i}),e_{m}\rangle.$ (6)
@@ 3.3 The Residual as Geometric Distance
$\|\phi(x_{i})\|^{2}=\|\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|^{2}+\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|^{2}.$ (7)
The second term is precisely the squared distance of the point $\phi(x_{i})$ from the subspace $S_{m-1}$. Since $\|\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|^{2}=\sum_{j=1}^{m-1}\langle\phi(x_{i}),e_{j}\rangle^{2}=\sum_{j=1}^{m-1}L_{i,j}^{2}$, the value stored in $d[i]$ at step $m$ corresponds to this residual energy:
$d[i]=\|\phi(x_{i})\|^{2}-\sum_{j=1}^{m-1}L_{i,j}^{2}=\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|^{2}.$ (8)
$S_{m-1}$$\|\phi(x_{i})\|$$\phi(x_{i})$$\text{Proj}_{S_{m-1}}(\phi(x_{i}))$$\sqrt{d[i]}$(Residual)
@@ 3.5 Formal Equivalence
We proceed by induction. For $m=1$, the subspace $S_{0}$ is the zero vector. Thus, $d[i]=k(x_{i},x_{i})=\langle\phi(x_{i}),\phi(x_{i})\rangle=\|\phi(x_{i})\|^{2}$, which is the squared distance to the zero vector. Assume the hypothesis holds for step $m$. The update rule for the diagonal is $d^{(new)}[i]=d^{(old)}[i]-L_{i,m}^{2}$. Using the inductive hypothesis and Equation (7), we have:
$d^{(new)}[i]=\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|^{2}-|\langle\phi(x_{i}),e_{m}\rangle|^{2}.$ (11)
Since $e_{m}$ is orthogonal to $S_{m-1}$ and $\|e_{m}\|=1$, removing the component projected onto $e_{m}$ corresponds to computing the squared distance to the larger subspace $S_{m}$, which is spanned by adding $e_{m}$ to $S_{m-1}$ (formally, $S_{m}=S_{m-1}\oplus\text{span}\{e_{m}\}$). ∎

### v3

@@ 3.2 The Update Step as Orthogonal Projection
Note that by definition of orthogonal projection, the residual vector $r_{m}$ is orthogonal to the subspace $S_{m-1}$. Since the diagonal element serves as the normalization factor $L_{mm}=\sqrt{d[\pi_{m}]}=\|r_{m}\|_{\mathcal{H}}$, the update step explicitly computes the projection onto the new normalized basis vector $e_{m}=r_{m}/\|r_{m}\|_{\mathcal{H}}$:
$L_{i,m}=\frac{\langle\phi(x_{i}),r_{m}\rangle_{\mathcal{H}}}{\|r_{m}\|_{\mathcal{H}}}=\langle\phi(x_{i}),e_{m}\rangle_{\mathcal{H}}.$ (6)
@@ 3.3 The Residual as Geometric Distance
$\|\phi(x_{i})\|_{\mathcal{H}}^{2}=\|\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|_{\mathcal{H}}^{2}+\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|_{\mathcal{H}}^{2}.$ (7)
The second term is precisely the squared distance of the point $\phi(x_{i})$ from the subspace $S_{m-1}$. Since $\|\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|_{\mathcal{H}}^{2}=\sum_{j=1}^{m-1}\langle\phi(x_{i}),e_{j}\rangle_{\mathcal{H}}^{2}=\sum_{j=1}^{m-1}L_{i,j}^{2}$, the value stored in $d[i]$ at step $m$ corresponds to this residual energy:
$d[i]=\|\phi(x_{i})\|_{\mathcal{H}}^{2}-\sum_{j=1}^{m-1}L_{i,j}^{2}=\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|_{\mathcal{H}}^{2}.$ (8)
$S_{m-1}$$\|\phi(x_{i})\|_{\mathcal{H}}$$\phi(x_{i})$$\text{Proj}_{S_{m-1}}(\phi(x_{i}))$$\sqrt{d[i]}$(Residual)
@@ 3.5 Formal Equivalence
We proceed by induction. For $m=1$, the subspace $S_{0}$ is the zero vector. Thus, $d[i]=k(x_{i},x_{i})=\langle\phi(x_{i}),\phi(x_{i})\rangle_{\mathcal{H}}=\|\phi(x_{i})\|_{\mathcal{H}}^{2}$, which is the squared distance to the zero vector. Assume the hypothesis holds for step $m$. The update rule for the diagonal is $d^{(new)}[i]=d^{(old)}[i]-L_{i,m}^{2}$. Using the inductive hypothesis and Equation (7), we have:
$d^{(new)}[i]=\|\phi(x_{i})-\text{Proj}_{S_{m-1}}(\phi(x_{i}))\|_{\mathcal{H}}^{2}-|\langle\phi(x_{i}),e_{m}\rangle_{\mathcal{H}}|^{2}.$ (11)
Since $e_{m}$ is orthogonal to $S_{m-1}$ and $\|e_{m}\|_{\mathcal{H}}=1$, removing the component projected onto $e_{m}$ corresponds to computing the squared distance to the larger subspace $S_{m}$, which is spanned by adding $e_{m}$ to $S_{m-1}$ (formally, $S_{m}=S_{m-1}\oplus\text{span}\{e_{m}\}$). ∎
@@ 5 Discussion
  • Point-wise vs. Subspace FPS: It is crucial to distinguish between two common interpretations of "Farthest Point Sampling." Standard geometric FPS typically selects the point maximizing the distance to the nearest neighbor in the selected set ($x_{next}=\arg\max_{x}\min_{s\in S}\|x-s\|$). Even if implemented in the feature space using the kernel metric ($d(x,s)=\|\phi(x)-\phi(s)\|_{\mathcal{H}}$), this approach fundamentally places spheres around selected points. In contrast, Pivoted Cholesky acts as a Subspace FPS in the RKHS, selecting the point maximizing the orthogonal distance to the linear subspace spanned by the selected feature vectors ($x_{next}=\arg\max_{x}\|\phi(x)-\text{Proj}_{\text{span}(S)}(\phi(x))\|_{\mathcal{H}}$). This distinction is critical: a point can be far from any specific selected point (high pairwise distance) yet linearly dependent on their combination (zero Cholesky residual). Thus, Cholesky maximizes the "volume" spanned by the features rather than merely filling the space. Note, however, that for stationary kernels with very narrow bandwidths (fast decay), the feature vectors become nearly orthogonal. In this specific regime, the distance to the subspace is dominated by the distance to the nearest neighbor, and the two sampling strategies effectively coincide.